OpenStudy (anonymous):

Find the indicated power using De Moivre's Theorem. (Express your fully simplified answer in the form a + bi.) (1 − i )^2

OpenStudy (bradely):

(1-i)^2 (a-b)^2 =a^2-2ab+b^2 1^2-(2i)+i^2 1-2i-1 since i^2=-1 0-2i Source: http://www.mathskey.com/question2answer/